# MToP: A MATLAB Benchmarking Platform for Evolutionary Multitasking

### v5

@@ 1. Introduction
Evolutionary computation (EC), inspired by natural evolution, has experienced rapid growth owing to its effectiveness and efficiency. Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and ease of implementation. Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization (Wang2020CORCO; Tian2021CCMO), multi-objective optimization (Zhang2007MOEA-D; Deb2014NSGA3), and combinatorial optimization (Feng2021VRPHTO; Feng2021EEMTA). In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking (Wei2021Review-MTO). A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ minimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
@@
where $\Omega_{k}$ and $\bm{F}_{k}$ are the decision space and objective function of the $k$-th task. Note that for a multi-objective multi-task optimization problem, the $\bm{F}_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of Pareto sets for the $k$-th task. For task $k$ with upper and lower bounds $\bm{L}_{k}$ and $\bm{U}_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space (Gupta2016MFEA) as follows:
@@
In addition, the dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$. During objective function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision. EMT has found successful applications in various domains, including engineering scheduling (Feng2021VRPHTO; Gupta2021Half), nonlinear equation systems (Li2024Evolutionary), feature selection in machine learning (Chen2021Evolutionary), anomaly detection (Wang2022AUC), point cloud registration (Wu2024Evolutionary; Wu2023Evolutionary), and reinforcement learning (Zhang2023Multitask).
@@
To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques (Tan2023Knowledge; Gupta2018Insights). Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources (Tan2021ETO; Li2024MFEA-RL). The first attempt of EMT can be traced back to the multifactorial EA (Gupta2016MFEA), which introduced an implicit knowledge representation via random mating among optimization tasks. Subsequently, numerous EAs tailored for MTO have emerged. These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks (Bali2020MFEA2; Zhou2021MFEA-AKT), or a multi-population framework allocating separate populations for each task (Li2020MFMP; Li2023MTEA-SaO; Li2024MTES-KG). Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space (Gupta2016MFEA; Zhou2021MFEA-AKT), auto-encoding (Feng2019EMEA; Zhou2021Learnable; Gu2025MTEA-PAE), affine transformation (Xue2022AT-MFEA; Lin2024Ensemble), and adversarial generative models (Liang2023EMT-GS) have been proposed in MTEAs. Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment (Bali2020MFEA2; Jiang2023BoKT), knowledge selection (Wang2022MTEA-AD; Liang2022EMaTO-MKT; Li2025MTDE-MKTA), and historical feedback (Li2022CompetitiveMTO; Lin2021EMT-ET; Li2023MTSRA), have also been investigated. Despite the proliferation of MTEAs proposed by researchers, there is currently no standardized programming language, code pattern, or software platform for EMT source codes. This presents challenges for newcomers entering the field of EMT and for researchers seeking to conduct convenient experimental comparisons of algorithms.
@@
To overcome these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field. MToP is designed to provide a comprehensive software platform for researchers to evaluate MTEAs on benchmark MTO problems and explore real-world applications. The selection of MATLAB is based on several strategic factors. First, many foundational and seminal works in the EMT field, including the original MFEA (Gupta2016MFEA) and MFEA-II (Bali2020MFEA2), were developed and open-sourced in MATLAB. This established a strong precedent and a rich ecosystem of existing code for the community. Second, the prominence of other successful EC platforms in MATLAB, most notably PlatEMO (Tian2017PlatEMO), provides a high-quality open-source reference. This synergy allows for the adaptaion and reuse of well-vetted modules, such as those for multi-objective optimization, which is crucial for the multi-objective MTO subfield. Finally, MATLAB’s powerful numerical computing environment and its capabilities for building user-friendly graphical interfaces make it an ideal choice for a comprehensive platform. The main contributions of MToP are summarized as follows:
@@ 2. Related Work
  • PlatEMO (Tian2017PlatEMO): A comprehensive MATLAB platform tailored for evolutionary multi-objective optimization.
  • EDOLAB (Peng2023EDOLAB): A MATLAB-based platform focusing on evolutionary dynamic optimization.
  • IOHprofiler (Doerr2018IOHprofiler): A modular framework, including IOHexperimenter (deNobel2024IOHexperimenter) and IOHanalyzer (Wang2022IOHanalyzer), for the detailed benchmarking and analysis of iterative optimization heuristics, with a primary focus on single-objective and multi-objective problems.
  • EvoX (Huang2024EvoX): A distributed GPU-accelerated library focused on scalability and expediting complex optimization and reinforcement learning tasks.
  • PyPop7 (Duan2024PyPop7): A pure-Python library for population-based single-objective black-box optimization, particularly for large-scale optimization.
  • MetaBox (Ma2023MetaBox): A Python-based platform for meta-black-box optimization, which uses meta-learning to design optimizers, with its latest version using MTO as a training scenario.
@@
PlatEMO (Tian2017PlatEMO) MATLAB Evolutionary Multi-objective Optimization Partial (v4)
EDOLAB (Peng2023EDOLAB) MATLAB Evolutionary Dynamic Optimization No
IOHprofiler (Doerr2018IOHprofiler) C++/Python/R Benchmarking Iterative Optimization Heuristics No
EvoX (Huang2024EvoX) Python Distributed GPU-accelerated Optimization No
PyPop7 (Duan2024PyPop7) Python Large-scale Black-box Optimization No
MetaBox (Ma2023MetaBox) Python Meta-black-box Optimization via Meta-learning As a test scenario (v2)
@@
A review of these tools, summarized in Table 1, justifies the development of a new, dedicated framework. At the time MToP development commenced [^1] MToP has been continuously updated since Sep. 21, 2021 at: https://github.com/intLyc/MTO-Platform, a platform with native, comprehensive support for MTO benchmarking was absent. The most prominent existing platform, PlatEMO (Tian2017PlatEMO), has a core architecture that is multi-objective optimization-centric, meticulously designed for solving single-task multi-objective optimization. This presents a fundamental architectural mismatch for EMT. Unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks, the implementation of inter-task solution space mapping, and diverse knowledge transfer techniques. Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics, which brings uncertainty to the pattern of results display and analysis. These distinctive requirements present significant challenges for implementation in platforms not designed for them. Consequently, extending current framework like PlatEMO to natively handle these MTO-specific concepts is non-trivial and leads to significant extensibility limitations. This is particularly evident in the context of many-task optimization, where PlatEMO’s support for specialized multitask metrics is restricted. Although preliminary MTO support was added to PlatEMO (v4.0) [^2] PlatEMO v4.0 (released on Oct. 13, 2022) introduced preliminary MTO support., these additions are supplementary rather than architecturally native. The other platforms listed in Table 1 are similarly specialized for different, non-MTO-centric goals. Moreover, EMT has been extended to the subfields of many-task optimization (Chen2020MaTDE; Liaw2019SBO; Li2024TNG-NES), multi-objective MTO (Gupta2017MO-MFEA; Li2025MTEA-DCK), competitive MTO (Li2022CompetitiveMTO; Li2023MTSRA; Li2024CMO-MTO), and constrained multi-task optimization (Li2022CMTO-Benchmark; Zhang2024CEDA). This growing complexity further necessitates a dedicated, MTO-native framework. While a new framework was deemed necessary, MToP actively integrates and appropriately cites code modules from these established platforms where feasible, thereby reducing redundant development and focusing on its core MTO-specific contributions.
@@ 3.1.2. Experiment Module
The module supports saving the complete, raw experimental dataset (MTOData) to a .mat file, which can be fully reloaded into the platform at any time. Once the data is generated, the module offers a suite of tools for post-processing and analysis. This includes functions to automatically calculate performance metrics, which are categorized to support both single-task and multi-task indicators, and to perform statistical significance tests to compare algorithmic performance. All generated tables and test results can be exported to .xlsx, .csv, or .tex formats. Furthermore, it provides granular export options for specific use cases, such as exporting convergence data in .csv format compatible with IOHanalyzer (Wang2022IOHanalyzer) or saving the optimal solution (single-objective MTO) or set of Pareto sets (multi-objective MTO) to .mat files for external study.
Finally, the module provides essential visualization capabilities for interpreting the experimental outcomes. As shown in Fig. 2, this includes generating convergence plots to simultaneously compare the performance trajectories of multiple algorithms across various problems (Fig. 2 (a)), a function applicable to both single- and multi-objective problems. Furthermore, for multi-objective MTO tasks, it can plot the acquired Pareto fronts (Fig. 2 (b)) to visually assess the quality of the final solution sets. Since all visualizations are generated as standard MATLAB figures, users can leverage MATLAB’s built-in plotting tools for further customization, annotation, and export.
@@ 3.2.1. Class Diagram
  • Function module groups static utility functions such as metrics (e.g., IGD()) and evolutionary operators (e.g., GA_Crossover()).
@@ 3.2.2. Sequence Workflow
Figure 4. Sequence diagram of MToP. The main workflow involves MTO_GUI / MTO_CMD, Algorithm, Problem, and Metric.
Fig. 4 depicts the dynamic interaction workflow between the main components when running an experiment:
@@
  • Then, the interface handles the setup phase, configuring the Algorithm and Problem instances by calling setParameter() and setTasks(). Next, the interface invokes the run() method of the Algorithm. The Algorithm performs its Initialization() and then enters the main optimization loop, conditioned on while notTerminated().
  • Inside the loop, the Algorithm executes its Evaluation() method, which subsequently calls the evaluate(individual) method on the Problem object. The Problem returns the calculated Obj, Con values. Following evaluation, the Algorithm calls EvolvePopulation() to generate the next population.
  • Once the loop terminates, the Algorithm returns the bestResult to the GUI/CMD interface. The interface then passes this result to the Metric module for Calculation(). Finally, the computed metricValue is returned and displayed to the user.
@@ 3.2.3. File Structure
  • Metrics/: it contains all result evaluation metrics, organized into subfolders for single-objective Metrics/Single-objective/ and multi-objective Metrics/Multi-objective optimization. Specific metric files include Obj.m, Obj_MTS.m, and IGD.m, among others.
@@ 3.2.4. Experiment Data Structure
  • Results: This is the $P\times A\times R$ struct array containing the core optimization output. Each struct MTOData.Results(i, j, k) contains the detailed results for a single run, aggregated across all $T$ tasks and all $G$ saved convergence points (where $G$ is the Results Num). Each struct contains:
    • Obj: Stores the objective values. For single-objective problems, this is a $T\times G$ numerical array storing the single best value for each task at each convergence point. For multi-objective problems, this is a $T\times 1$ cell array, where each cell Obj{t} contains a $G\times N\times M_{t}$ array representing the full objective values of $N$ solutions.
@@ 3.3.2. Problem
Moreover, for multi-objective optimization, the external function getOptimum() is provided. It defines the true optimal solutions (i.e., true Pareto front) if known, which is required for metrics like IGD. For problems where the true front is unknown (e.g., real-world applications), this function is instead used to provide the reference point for calculating metrics such as hypervolume (HV).

### v6

@@ 1. Introduction
Evolutionary computation (EC), inspired by natural evolution, has experienced rapid growth owing to its effectiveness and efficiency. Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and ease of implementation. Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization (Wang et al., 2020; Tian et al., 2021), multi-objective optimization (Zhang and Li, 2007; Deb and Jain, 2014), and combinatorial optimization (Feng et al., 2021b, a). In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking (Wei et al., 2022). A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ minimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
@@
where $\Omega_{k}$ and $\bm{F}_{k}$ are the decision space and objective function of the $k$-th task. Note that for a multi-objective multi-task optimization problem, the $\bm{F}_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of Pareto sets for the $k$-th task. For task $k$ with upper and lower bounds $\bm{L}_{k}$ and $\bm{U}_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space (Gupta et al., 2016) as follows:
@@
In addition, the dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$. During objective function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision. EMT has found successful applications in various domains, including engineering scheduling (Feng et al., 2021b; Gupta et al., 2022), nonlinear equation systems (Li et al., 2024a), feature selection in machine learning (Chen et al., 2022), anomaly detection (Wang et al., 2022b), point cloud registration (Wu et al., 2024a, b), and reinforcement learning (Zhang et al., 2023).
@@
To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques (Tan et al., 2023; Gupta et al., 2018). Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources (Tan et al., 2021; Li et al., 2024b). The first attempt of EMT can be traced back to the multifactorial EA (Gupta et al., 2016), which introduced an implicit knowledge representation via random mating among optimization tasks. Subsequently, numerous EAs tailored for MTO have emerged. These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks (Bali et al., 2020; Zhou et al., 2021b), or a multi-population framework allocating separate populations for each task (Li et al., 2020, 2023b, 2024d). Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space (Gupta et al., 2016; Zhou et al., 2021b), auto-encoding (Feng et al., 2019; Zhou et al., 2021a; Gu et al., 2025), affine transformation (Xue et al., 2022; Lin et al., 2024), and adversarial generative models (Liang et al., 2023) have been proposed in MTEAs. Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment (Bali et al., 2020; Jiang et al., 2023), knowledge selection (Wang et al., 2022a; Liang et al., 2022; Li and Gong, 2025), and historical feedback (Li et al., 2022a; Lin et al., 2021; Li et al., 2023a), have also been investigated. Despite the proliferation of MTEAs proposed by researchers, there is currently no standardized programming language, code pattern, or software platform for EMT source codes. This presents challenges for newcomers entering the field of EMT and for researchers seeking to conduct convenient experimental comparisons of algorithms.
@@
To overcome these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field. MToP is designed to provide a comprehensive software platform for researchers to evaluate MTEAs on benchmark MTO problems and explore real-world applications. The selection of MATLAB is based on several strategic factors. First, many foundational and seminal works in the EMT field, including the original MFEA (Gupta et al., 2016) and MFEA-II (Bali et al., 2020), were developed and open-sourced in MATLAB. This established a strong precedent and a rich ecosystem of existing code for the community. Second, the prominence of other successful EC platforms in MATLAB, most notably PlatEMO (Tian et al., 2017), provides a high-quality open-source reference. This synergy allows for the adaptaion and reuse of well-vetted modules, such as those for multi-objective optimization, which is crucial for the multi-objective MTO subfield. Finally, MATLAB’s powerful numerical computing environment and its capabilities for building user-friendly graphical interfaces make it an ideal choice for a comprehensive platform. The main contributions of MToP are summarized as follows:
@@ 2. Related Work
  • PlatEMO (Tian et al., 2017): A comprehensive MATLAB platform tailored for evolutionary multi-objective optimization.
  • EDOLAB (Peng et al., 2023): A MATLAB-based platform focusing on evolutionary dynamic optimization.
  • IOHprofiler (Doerr et al., 2018): A modular framework, including IOHexperimenter (de Nobel et al., 2024) and IOHanalyzer (Wang et al., 2022c), for the detailed benchmarking and analysis of iterative optimization heuristics, with a primary focus on single-objective and multi-objective problems.
  • EvoX (Huang et al., 2024): A distributed GPU-accelerated library focused on scalability and expediting complex optimization and reinforcement learning tasks.
  • PyPop7 (Duan et al., 2024): A pure-Python library for population-based single-objective black-box optimization, particularly for large-scale optimization.
  • MetaBox (Ma et al., 2023): A Python-based platform for meta-black-box optimization, which uses meta-learning to design optimizers, with its latest version using MTO as a training scenario.
@@
PlatEMO (Tian et al., 2017) MATLAB Evolutionary Multi-objective Optimization Partial (v4)
EDOLAB (Peng et al., 2023) MATLAB Evolutionary Dynamic Optimization No
IOHprofiler (Doerr et al., 2018) C++/Python/R Benchmarking Iterative Optimization Heuristics No
EvoX (Huang et al., 2024) Python Distributed GPU-accelerated Optimization No
PyPop7 (Duan et al., 2024) Python Large-scale Black-box Optimization No
MetaBox (Ma et al., 2023) Python Meta-black-box Optimization via Meta-learning As a test scenario (v2)
@@
A review of these tools, summarized in Table 1, justifies the development of a new, dedicated framework. At the time MToP development commenced [^1] MToP has been continuously updated since Sep. 21, 2021 at: https://github.com/intLyc/MTO-Platform, a platform with native, comprehensive support for MTO benchmarking was absent. The most prominent existing platform, PlatEMO (Tian et al., 2017), has a core architecture that is multi-objective optimization-centric, meticulously designed for solving single-task multi-objective optimization. This presents a fundamental architectural mismatch for EMT. Unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks, the implementation of inter-task solution space mapping, and diverse knowledge transfer techniques. Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics, which brings uncertainty to the pattern of results display and analysis. These distinctive requirements present significant challenges for implementation in platforms not designed for them. Consequently, extending current framework like PlatEMO to natively handle these MTO-specific concepts is non-trivial and leads to significant extensibility limitations. This is particularly evident in the context of many-task optimization, where PlatEMO’s support for specialized multitask metrics is restricted. Although preliminary MTO support was added to PlatEMO (v4.0) [^2] PlatEMO v4.0 (released on Oct. 13, 2022) introduced preliminary MTO support., these additions are supplementary rather than architecturally native. The other platforms listed in Table 1 are similarly specialized for different, non-MTO-centric goals. Moreover, EMT has been extended to the subfields of many-task optimization (Chen et al., 2020; Liaw and Ting, 2019; Li et al., 2024c), multi-objective MTO (Gupta et al., 2017; Li et al., 2025), competitive MTO (Li et al., 2022a, 2023a, 2024e), and constrained multi-task optimization (Li et al., 2022b; Zhang et al., 2024). This growing complexity further necessitates a dedicated, MTO-native framework. While a new framework was deemed necessary, MToP actively integrates and appropriately cites code modules from these established platforms where feasible, thereby reducing redundant development and focusing on its core MTO-specific contributions.
@@ 3.1. Functional Modules
Notably, MToP is architecturally designed without any inherent software limitation on the number of tasks. Users can programmatically define benchmarks with an arbitrary number of tasks without hard-coded limits.
@@ 3.1.2. Experiment Module
The module supports saving the complete, raw experimental dataset (MTOData) to a .mat file, which can be fully reloaded into the platform at any time. Once the data is generated, the module offers a suite of tools for post-processing and analysis. This includes functions to automatically calculate performance metrics, which are categorized to support both single-task and multi-task indicators, and to perform statistical significance tests to compare algorithmic performance. All generated tables and test results can be exported to .xlsx, .csv, or .tex formats. Furthermore, it provides granular export options for specific use cases, such as exporting convergence data in .csv format compatible with IOHanalyzer (Wang et al., 2022c) or saving the optimal solution (single-objective MTO) or set of Pareto sets (multi-objective MTO) to .mat files for external study.
Finally, the module provides essential visualization capabilities for interpreting the experimental outcomes. As shown in Fig. 2, this includes generating convergence plots to simultaneously compare the performance trajectories of multiple algorithms across various problems (Fig. 2 (a)), a function applicable to both single- and multi-objective problems. Furthermore, for multi-objective MTO tasks, it can plot the acquired Pareto front approximations (Fig. 2 (b)) to visually assess the quality of the final solution sets. Since all visualizations are generated as standard MATLAB figures, users can leverage MATLAB’s built-in plotting tools for further customization, annotation, and export.
@@ 3.2.1. Class Diagram
  • Function module groups static utility functions such as metrics (e.g., IGDp(), i.e., IGD+) and evolutionary operators (e.g., GA_Crossover()).
@@ 3.2.2. Sequence Workflow
Figure 4. Sequence diagram of MToP. The main workflow involves MTO_GUI / MTO_CMD, Algorithm (MFEA), Problem (CMT1), and Metric (Obj).
Fig. 4 illustrates the dynamic interaction workflow using a concrete example of the MFEA algorithm solving the CMT1 problem:
@@
  • The interface configures the specific algorithm (e.g., MFEA) and problem (e.g., CMT1) instances by calling setParameter() and setTasks(). Subsequently, the interface invokes the run() method of MFEA, which performs Initialization() and enters the main optimization loop conditioned on while notTerminated().
  • Inside the loop, the workflow first executes the Generation() phase. Here, the algorithm calls specific operators, such as GA_Crossover() and GA_Mutation(), to produce offspring.
  • Following generation, the Evaluation() method is triggered. This invokes the evaluate() method on the CMT1 object, which returns the objective and constraint values (Obj, Con). Finally, the Selection() strategy is applied to determine the survivors for the next generation.
  • Once the loop terminates, MFEA returns the bestResult to the interface. The interface then passes this result to the metric module (e.g., Obj) for Calculation(). Finally, the computed metricValue is displayed to the user.
@@ 3.2.3. File Structure
  • Metrics/: it contains all result evaluation metrics, organized into subfolders for single-objective Metrics/Single-objective/ and multi-objective Metrics/Multi-objective optimization. Specific metric files include Obj.m, Obj_MTS.m, and IGDp.m, among others.
@@ 3.2.4. Experiment Data Structure
  • Results: This is the $P\times A\times R$ struct array containing the core optimization output. Each struct MTOData.Results(i, j, k) contains the detailed results for a single run, aggregated across all $T$ tasks and all $G$ saved checkpoints (where $G$ is the Results Num). Each struct contains:
    • Obj: Stores the objective values. For single-objective problems, this is a $T\times G$ numerical array storing the single best value for each task at each checkpoint. For multi-objective problems, this is a $T\times 1$ cell array, where each cell Obj{t} contains a $G\times N\times M_{t}$ array representing the full objective values of $N$ solutions.
@@ 3.3.2. Problem
Moreover, for multi-objective optimization, the external function getOptimum() is provided. It defines the true optimal solutions (i.e., true Pareto front) if known, which is required for metrics like IGD+. For problems where the true front is unknown (e.g., real-world applications), this function is instead used to provide the reference point for calculating metrics such as hypervolume (HV).